OpenStudy (anonymous):

MEDAL AND FAN AWAREDED!!!!! Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3. y - 6 = -3x - 3 y - 6 = -3(x + 1) y = -3x + 3 3x + y = 3

Directrix (directrix):

The general form looks something like this: y - y1 = m ( x - x1) where m is the slope of the line and (x1, y1) is a point on the line.

Directrix (directrix):

(-1, 6) is your (x1, y1) and your m = -3 So, see if you can fit that information into the general point-slope equation.

OpenStudy (anonymous):

can you show me how to do that?

Directrix (directrix):

I will if you will try first. @CoinBank77

OpenStudy (anonymous):

alright -3 = (-1 - 6) is it something like that?

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Not quite. y - y1 = m ( x - x1)

Directrix (directrix):

y - y1 = m ( x - x1) m = -3

Directrix (directrix):

y - y1 = -3 * ( x - x1) @CoinBank77 Into the above equation, see if you can fit the point (-1, 6). x1 = -1 and y1 = 6

Directrix (directrix):

Give it a try, okay?

OpenStudy (anonymous):

Alright, y - y1 = -3 ( x - 6) am i off in the right direction?

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Yes, you are.

OpenStudy (anonymous):

y - 6 = -3 ( x - 6)?

Directrix (directrix):

Just a second.

Directrix (directrix):

No. I led you astray. Sorry. I mixed the x1 with the y1. y - y1 = -3 ( x - x1) Think of it this way: y - 6 = -3 ( x - (-1) )

Directrix (directrix):

What is the negative of -1 ? We have to clean up this: - (-1)

Directrix (directrix):

y - 6 = -3 ( x - (-1) ) y - 6 = -3 ( x + 1) ---> Final Do you agree?

OpenStudy (anonymous):

Negative of -1 is a positive 1

Directrix (directrix):

@CoinBank77 Of your answer options, which one is correct? Do you see it.

OpenStudy (anonymous):

Well from what you've helped me with, it should be: y - 6 = -3(x + 1) I filled in the -3 for m, 1 for y, so that's the answer?

Directrix (directrix):

We have to leave this in point-slope form so to the best of my thinking, we stop with this option: y - 6 = -3(x + 1) which I think is Option B. Again, I apologize for getting the x1 mixed with the y1 in my head and then passing that to you.

OpenStudy (anonymous):

It's all good, I just submitted the question. It is B!! Thank you for spending all this time with me, I appreciate it :)!

OpenStudy (anonymous):

It would've taken me a much longer time trying to figure it out on my own haha.

Directrix (directrix):

You are welcome.